a 95% confidence interval for the mean of a population is given as (3.0, 5.0). If a sample size of 100 was used to obtain this interval what was the standard deviation used in the calculation?
Answer by stanbon(75887)

a 95% confidence interval for the mean of a population is given as (3.0, 5.0). If a sample size of 100 was used to obtain this interval what was the standard deviation used in the calculation?
-----------------
2*ME = 5-3 = 2
ME = 1
---------------
So zs/sqrt(n) = 1
1.96s/sqrt(100) = 0.196s = 1
s = 5.1020
==================
Cheers,
Stan H.